OpenStudy (anonymous):

Write the equation of the line in slope-intercept form given the point (4,3) and the slope of 1/2.

OpenStudy (anonymous):

if you are given a point and a slope you can put it in point slope form: y-y1 = m(x-x1) where m is the slope and (x1, y1) is the given point.

OpenStudy (anonymous):

since m=1/2 and (x1,y1) is (4,3), then the point-slope form is y-3=(1/2)(x-4).

OpenStudy (anonymous):

since you want this in slope intercept form multiply out the right side and bring over the 3 from the left side to the right. then simplify.